OpenStudy (anonymous):

using f(x) = 3x - 3 and g(x) = 2x, what is f(4) - g(2) ? a. 5 b. -5 c. 10 d. -10 **May someone explain how to work this problem..

OpenStudy (jojo4eva):

a. 5

OpenStudy (jojo4eva):

3(4)-3=12-3=9 2(2)=4 9-4=5 which gives you a final answer of 5

OpenStudy (solomonzelman):

\(\Large\color{black}{ f(4)}\) means that you plug in \(\Large\color{black}{ 4}\) into f(x) for x. For example, f(2) \(\Large\color{black}{ f(2)=3(2)-3~~~~=6-3~~~~=3}\) So can you find f(4), (the same way) ? Similarly, g(2) means that you plug in 2 into g(x) .... so, for example, g(5) \(\Large\color{black}{ g(5) = 2(5)~~~=10}\) Can you find g(2) ?

OpenStudy (solomonzelman):

After you find f(4) and g(2), substitute the found values, into "f(4) - g(2)"